Mn) xin) -In 0.5 0.5 (2-m)s I-m)s f4-m)s 3-ms PN 01 2 3 4 2(n) 0.5

How to get Z(n)? especially n=2,n=3.
Answer #1

Comoldiob Z(n)= 5 x(K) YIn-K) K 42 Z(2) x(K) g(2) K o + (o-t)h (0) K = 2) (2-2)>X + = (0) () + from grabh 2 z(3) XKy(3 -k). KPlease let me know if you have any doubt or if you found any discrepancy. Thanks u too :)